OpenStudy (anonymous):

evaluate the expression for the given values. If A =5 and B= -5 , find a-b^2

OpenStudy (texaschic101):

we are basically just subbing and solving.. a = 5 and b = -5 a - b^2 = 5 - (-5^2) = 5 - 25 = -20 <== your answer any questions ?

OpenStudy (anonymous):

lots lol but i wont bother